# Fine-Tuning on Noisy Instructions: Effects on Generalization and Performance

### v1

@@ 1 Introduction
Instruction-tuning is widely adopted to enable LLMs to follow complex instructions and respond properly (Sanh et al., 2022; Zhao et al., 2023; Chang et al., 2023; Minaee et al., 2024; Zhang et al., 2024). During instruction-tuning, LLMs are fine-tuned on datasets comprising various task instructions and their corresponding responses.
@@
LLMs have been shown to be sensitive to prompt variability, producing inconsistent responses when given semantically equivalent prompts Sun et al. (2024); Zhao et al. (2024); Yan et al. (2024). To remedy this, recent instruction datasets are often generated with extensive paraphrasing using LLMs to increase data diversity (Peng et al., 2023). However, this paraphrased data is of high quality with minimal noise in the instructions. A different line of work has explored the robustness of instruction-tuned models to instruction variations during inference by introducing different types of noise, such as deleting words (Gu et al., 2023). However, how noisy data may affect LLMs during training has yet to be explored.
In this paper, we focus on answering the following research question: Can fine-tuning of base models on perturbed instructions improve their resilience to noisy user inputs? Our question is theoretically motivated by previous work that has shown that introducing noise during training acts as a form of regularization (Bishop, 1995), which can prevent overfitting and improve generalization. [^2] We may interchangeably use terms such as ‘resilience’ or ‘robustness’ to refer to the model’s capacity to withstand and adapt to noisy user inputs without a substantial drop in performance. The term ‘generalization’ refers to the model’s capacity to perform a given task when presented with novel, previously unseen phrasings or formats.
To evaluate the impact of instruction perturbation and simulate noisy user inputs, we employ five strategies inspired by Gu et al. (2023): (1) delete stop words, (2) shuffle words, (3) delete words, (4) replace words, and (5) insert words. We further introduce a sixth perturbation strategy by adding misspellings. These strategies allow us to simulate and analyze various forms of perturbed instructions, including both structural and semantic changes. We construct four perturbed instruction-tuning datasets where GPT4-Alpaca (Peng et al., 2023), Super-Natural (Wang et al., 2022), and Dolly (Conover et al., 2023) are combined and then perturbed. We include versions of the data where 25%, 50%, 75%, and 100% of the instructions are perturbed respectively, which allows us to compare how LLMs are affected by different proportions of noisy instructions presented in instruction-tuning. We evaluate performance across three widely-used language understanding benchmarks: Massive Multitask Language Understanding (Hendrycks et al., 2021), Big-Bench Hard (Suzgun et al., 2022), and Grade School Math (Cobbe et al., 2021). Figure 1 shows the process of generating noisy instruction-tuning data and LLM fine-tuning.
@@ 2.1 Analyzing Instruction-tuning
Instruction fine-tuning enables LLMs to follow user instructions and reduces the need for few-shot in-context examples Ouyang et al. (2022); Wei et al. (2022a); Touvron et al. (2023a); Chung et al. (2024). The instruction-tuning datasets contain instructions of various tasks and their corresponding responses which can be human-annotated Mishra et al. (2022) or synthetically generated Taori et al. (2023); Peng et al. (2023).
AutoPrompt Shin et al. (2020), which applied a gradient-based search to optimize the prompt for various tasks, usually finds prompts that are hardly comprehensible by humans, indicating that language models have a vastly different way to understand instructions. Recent studies have investigated the internal mechanisms of instruction fine-tuning and their influence on LLMs. By observing the output token distribution shift of models before and after instruction-tuning, Lin et al. (2024) found that most shifts occur with stylistic tokens (e.g. discourse markers and transitional words), and knowledge content originates from untuned LLMs. By introducing knowledge interventions, Ren et al. (2024) also showed that instruction-tuning is a process of self-aligning the instructions with existing parametric knowledge rather than introducing new knowledge into the model.
@@ 2.2 Robustness of Instruction-tuned Models
Gu et al. (2023) investigated how instruction-tuned models handle instruction perturbations and paraphrasing. After fine-tuning a model on the original instructions training set and evaluating it on the perturbed instructions test set, they found the model was relatively robust in few-shot settings but notably sensitive in zero-shot scenarios. Similarly, Sun et al. (2024) showed that paraphrased instructions can disrupt model consistency and proposed a mitigation strategy using soft prompt embeddings to align semantically similar instructions. Wang et al. (2024) examined errors from speech recognition and OCR, finding such noise significantly degrades LLMs performance. They also explored using LLMs for zero-shot correction of noisy instructions. Yan et al. (2024) proposed contrastive instruction-tuning which align the hidden representations of instruction-instance pairs that are semantically equivalent but textually different and to differentiate those that are semantically distinct. Zhao et al. (2024) proposed a consistency alignment framework including instruction augmentation by paraphrasing and automatic self reward. Lou et al. (2024) introduced a dataset curation scheme by diversifying the task inputs with various facets. Kim et al. (2024) proposed instructive decoding, which improves instruction-following in instruction-tuned LLMs by contrasting the decoding process without additional tuning.
@@ 3 Instruction Perturbation Strategies
We investigate the impact of instruction fine-tuning on the performance of LLMs when subjected to noisy input conditions. Following Gu et al. (2023), we employ five instruction perturbation strategies: delete stop words, shuffle words, delete words, replace words, and insert words. Furthermore, we introduce misspelling as an additional noise injection approach. Table 1 shows an example instruction from the GPT4-Alpaca dataset (Peng et al., 2023). The model input consists of an instruction and associated context. The perturbation strategies are applied exclusively to the instruction component.
@@
We randomly select 25% of the words from an instruction and replace them using predictions from a pretrained BERT model (Devlin et al., 2019). Following Gu et al. (2023), we use BERT’s masked language modeling head to generate a contextually plausible substitute for all selected words. The selected words are replaced by [MASK] tokens, then BERT predicts replacement words in a forward pass. This strategy introduces minimal semantic shifts to the instructions without relying on a lexicon, which typically requires manual effort. This perturbation may or may not alter the core meaning of the instruction, depending on the replaced words and their context. The strategy allows us to investigate the model’s sensitivity to nuanced lexical variations and its ability to generalize under slightly altered task phrasing.
@@
We introduce additional words into the instruction by leveraging a pretrained BERT model (Devlin et al., 2019). Specifically, we randomly select positions between existing words, covering approximately 25% of the total word count, and insert a [MASK] token at each selected position. We then replace each [MASK] with words predicted by BERT, resulting in an augmented instruction containing additional, contextually plausible tokens. This perturbation may introduce noise, redundancy, or shifts in meaning, challenging the model’s ability to extract the core task intention from a more verbose or distorted input.
@@ 4.1 Models
We experiment with two open-weight base LLMs in two sizes: Qwen-2.5 (7B and 72B) (Yang et al., 2024); and Llama-3.1 (8B and 70B) (Dubey et al., 2024).
@@ 4.2 Instruction Datasets
GPT4-Alpaca (Peng et al., 2023) is derived from Alpaca (Taori et al., 2023), where the original examples are replaced with responses generated by GPT-4. Super-Natural Instruction (Wang et al., 2022) contains diverse tasks, including text classification and translation, with corresponding instructions. It is designed to evaluate the LLM abilities across a wide range of linguistic and functional contexts. Dolly (Conover et al., 2023) consists of instruction-following examples that reflect practical, real-world tasks like brainstorming and creative writing. The prompt-response pairs are high-quality and human-generated. Table 2 summarizes the number of samples in these datasets.
@@ 4.3 Implementation Details
We apply parameter-efficient fine-tuning methods for all experiments. Specifically, we use LoRA (Hu et al., 2022) to fine-tune the 7B and 8B models, and QLoRA (Dettmers et al., 2023) for the larger 70B and 72B models. Each model is fine-tuned for one epoch on each dataset mixture to ensure consistency across experiments. All fine-tuning runs were performed on a single NVIDIA H100 GPU. Full details on the fine-tuning hyperparameters are provided in Appendix A.
@@ 4.4 Evaluation
Massive Multitask Language Understanding (MMLU; Hendrycks et al. 2021):
@@
Big-Bench Hard (BBH; Suzgun et al. 2022):
A challenging subset of 23 tasks from the original BIG-Bench (Srivastava et al., 2023), aimed at evaluating advanced reasoning in language models. We assess both direct prompting and chain-of-thought (CoT) Wei et al. (2022b), using official prompts with three in-context examples, and report average exact match across sub-tasks.
Grade School Math (GSM8K; Cobbe et al. 2021):
@@
Additionally, to analyze potential side effects of instruction-tuning on noisy instructions, such as changes in toxicity or misinformation generation, we evaluate the models on: (1) ToxiGen (Hartvigsen et al., 2022) which measures the extent to which models generate toxic language and hate speech when explicitly prompted to do so across various demographic groups. We report the percentage of toxic outputs identified using a RoBERTa model (Liu et al., 2020) fine-tuned for toxicity detection, as described by Hartvigsen et al. (2022); and (2) TruthfulQA proposed by Lin et al. (2022) which assesses how effectively models can refrain from generating known falsehoods caused by misconceptions or false beliefs, while still generating informative and useful content. We use two off-the-shelf task-specific judge models developed by AllenAI based on Llama-2 (7B) (Touvron et al., 2023b) for measuring truthfulness [^3] https://huggingface.co/allenai/truthfulqa-truth-judge-llama2-7B and informativeness, following the setup of Groeneveld et al. (2024). In both datasets, we evaluate using only the original, unaltered prompts to measure model toxicity and truthfulness.
@@ 5 Results
Prior work has shown that CoT prompting outperforms direct prompting on benchmarks like BBH and GSM8K Kojima et al. (2022); Wei et al. (2022b), and our results confirm this, especially under instruction perturbation. On BBH (CoT), models like Llama-70B fine-tuned with fully perturbed instructions achieve the highest scores of 79.0% and 76.4% when evaluated with 25% and 50% perturbed instructions, respectively, indicating that CoT prompting benefits from increased robustness introduced during training. While Qwen-72B performs best when fine-tuned with less perturbed instructions, its relatively weaker performance under direct prompting suggests that incorporating perturbation during fine-tuning still contributes to improved generalization and reasoning robustness.
@@
We observe that the impact of instruction-tuning seems to vary by task. For example, MMLU shows relatively modest improvement from instruction-tuning, regardless of whether or not perturbations are applied. In contrast, BBH appears to consistently benefit from instruction-tuning, especially for Llama models. This observation aligns with findings from Sun and Dredze (2025), who suggest that certain tasks are already well-represented in a model’s pre-training data, leaving limited room for further gains through instruction-tuning. On the other hand, tasks that are underrepresented or poorly learned during pre-training can potentially see more substantial improvements as the model acquires new task-specific capabilities during instruction-tuning.
@@ 6.1 Safety and Bias
6.2 Individual Perturbation Strategies
6.3 Perturbation Intensity Ablation
@@ 7 Theoretical Grounding of Fine-Tuning on Noisy Instructions
Our findings suggest that incorporating perturbations during instruction-tuning may not only enhance model robustness to noisy or perturbed inputs but may also yield improvements on standard, unperturbed instructions. A plausible explanation for this effect is that noisy instruction-tuning serves as an implicit form of regularization Bishop (1995), potentially encouraging models to move beyond reliance on superficial linguistic patterns. Exposure to a wide spectrum of instruction formulations, including those containing syntactic or semantic anomalies, may discourage overfitting to narrow or canonical phrasing. These perturbations effectively broaden the training distribution, functioning as a form of data augmentation Dao et al. (2019); Hernández-García and König (2018); Vaibhav et al. (2019), and may thereby help LLMs to learn more robust and generalizable task representations. A particularly noteworthy observation is that models fine-tuned with 100% perturbed instructions often achieve high accuracy, even when evaluated on standard instructions. This may suggest that the perturbations could act not merely as noise, but as a potential source of useful inductive bias that enhances generalization across prompt formats.
@@ Acknowledgments
We would like to thank Atsuki Yamaguchi and the anonymous reviewers for their invaluable feedback. AA is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation grant EP/S023062/1.
@@ Appendix A Instruction-tuning Hyperparameters
For instruction-tuning and evaluation, we adopt the implementation from Open Instruct (Wang et al., 2023; Ivison et al., 2023, 2024; Lambert et al., 2024). Table 6 shows the hyperparameters used in our instruction fine-tuning experiments.

### v2

@@ 1 Introduction
Instruction-tuning is widely adopted to enable LLMs to follow complex instructions and respond properly (sanh2022multitask; zhao2023survey; Chang2023ASO; minaee2024large; zhang2024instructiontuninglargelanguage). During instruction-tuning, LLMs are fine-tuned on datasets comprising various task instructions and their corresponding responses.
@@
LLMs have been shown to be sensitive to prompt variability, producing inconsistent responses when given semantically equivalent prompts sun2024evaluating; zhao-etal-2024-improving; yan-etal-2024-contrastive. To remedy this, recent instruction datasets are often generated with extensive paraphrasing using LLMs to increase data diversity (peng2023instruction). However, this paraphrased data is of high quality with minimal noise in the instructions. A different line of work has explored the robustness of instruction-tuned models to instruction variations during inference by introducing different types of noise, such as deleting words (gu2023robustness). However, how noisy data may affect LLMs during training has yet to be explored.
In this paper, we focus on answering the following research question: Can fine-tuning of base models on perturbed instructions improve their resilience to noisy user inputs? Our question is theoretically motivated by previous work that has shown that introducing noise during training acts as a form of regularization (bishop1995training), which can prevent overfitting and improve generalization. [^2] We may interchangeably use terms such as ‘resilience’ or ‘robustness’ to refer to the model’s capacity to withstand and adapt to noisy user inputs without a substantial drop in performance. The term ‘generalization’ refers to the model’s capacity to perform a given task when presented with novel, previously unseen phrasings or formats.
To evaluate the impact of instruction perturbation and simulate noisy user inputs, we employ five strategies inspired by gu2023robustness: (1) delete stop words, (2) shuffle words, (3) delete words, (4) replace words, and (5) insert words. We further introduce a sixth perturbation strategy by adding misspellings. These strategies allow us to simulate and analyze various forms of perturbed instructions, including both structural and semantic changes. We construct four perturbed instruction-tuning datasets where GPT4-Alpaca (peng2023instruction), Super-Natural (wang2022super), and Dolly (conover2023free) are combined and then perturbed. We include versions of the data where 25%, 50%, 75%, and 100% of the instructions are perturbed respectively, which allows us to compare how LLMs are affected by different proportions of noisy instructions presented in instruction-tuning. We evaluate performance across three widely-used language understanding benchmarks: Massive Multitask Language Understanding (hendrycks2021measuring), Big-Bench Hard (Suzgun2022ChallengingBT), and Grade School Math (Cobbe2021TrainingVT). Figure 1 shows the process of generating noisy instruction-tuning data and LLM fine-tuning.
@@ 2.1 Analyzing Instruction-tuning
Instruction fine-tuning enables LLMs to follow user instructions and reduces the need for few-shot in-context examples ouyang2022training; wei2022finetuned; touvron2023llamaopenefficientfoundation; chung2024scaling. The instruction-tuning datasets contain instructions of various tasks and their corresponding responses which can be human-annotated mishra-etal-2022-cross or synthetically generated alpaca; peng2023instruction.
AutoPrompt shin-etal-2020-autoprompt, which applied a gradient-based search to optimize the prompt for various tasks, usually finds prompts that are hardly comprehensible by humans, indicating that language models have a vastly different way to understand instructions. Recent studies have investigated the internal mechanisms of instruction fine-tuning and their influence on LLMs. By observing the output token distribution shift of models before and after instruction-tuning, lin2024the found that most shifts occur with stylistic tokens (e.g. discourse markers and transitional words), and knowledge content originates from untuned LLMs. By introducing knowledge interventions, ren-etal-2024-learning also showed that instruction-tuning is a process of self-aligning the instructions with existing parametric knowledge rather than introducing new knowledge into the model.
@@ 2.2 Robustness of Instruction-tuned Models
gu2023robustness investigated how instruction-tuned models handle instruction perturbations and paraphrasing. After fine-tuning a model on the original instructions training set and evaluating it on the perturbed instructions test set, they found the model was relatively robust in few-shot settings but notably sensitive in zero-shot scenarios. Similarly, sun2024evaluating showed that paraphrased instructions can disrupt model consistency and proposed a mitigation strategy using soft prompt embeddings to align semantically similar instructions. wang-etal-2024-resilience examined errors from speech recognition and OCR, finding such noise significantly degrades LLMs performance. They also explored using LLMs for zero-shot correction of noisy instructions. In addition, abedin2025arithmattack demonstrated that LLMs are sensitive to noise in reasoning tasks by introducing random punctuation perturbations into math problem prompts.
yan-etal-2024-contrastive proposed contrastive instruction-tuning which align the hidden representations of instruction-instance pairs that are semantically equivalent but textually different while distinguishing those that are semantically different. zhao-etal-2024-improving proposed a consistency alignment framework that incorporates instruction augmentation through paraphrasing and automatic self-reward mechanisms. lou2024muffin introduced a dataset curation scheme that diversifies task inputs across multiple facets to enhance instruction diversity. kim2024instructive presented instructive decoding, a method that strengthens instruction-following abilities in instruction-tuned LLMs by contrasting decoding paths without additional fine-tuning.
luo2024robustft introduced a denoising framework that detects and filters out low-quality samples, enabling more robust fine-tuning of models on noisy downstream datasets. Similarly, agrawal2025enhancing performed an empirical analysis focused on enhancing model robustness to character- and word-level perturbations in classification tasks, revealing that iterative self-denoising surpasses approaches like ensembling and representation alignment.
@@ 3 Instruction Perturbation Strategies
We investigate the impact of instruction fine-tuning on the performance of LLMs when subjected to noisy input conditions. Following gu2023robustness, we employ five instruction perturbation strategies: delete stop words, shuffle words, delete words, replace words, and insert words. Furthermore, we introduce misspelling as an additional noise injection approach. Table 1 shows an example instruction from the GPT4-Alpaca dataset (peng2023instruction). The model input consists of an instruction and associated context. The perturbation strategies are applied exclusively to the instruction component.
@@
We randomly select 25% of the words from an instruction and replace them using predictions from a pretrained BERT model (devlin-etal-2019-bert). Following gu2023robustness, we use BERT’s masked language modeling head to generate a contextually plausible substitute for all selected words. The selected words are replaced by [MASK] tokens, then BERT predicts replacement words in a forward pass. This strategy introduces minimal semantic shifts to the instructions without relying on a lexicon, which typically requires manual effort. This perturbation may or may not alter the core meaning of the instruction, depending on the replaced words and their context. The strategy allows us to investigate the model’s sensitivity to nuanced lexical variations and its ability to generalize under slightly altered task phrasing.
@@
We introduce additional words into the instruction by leveraging a pretrained BERT model (devlin-etal-2019-bert). Specifically, we randomly select positions between existing words, covering approximately 25% of the total word count, and insert a [MASK] token at each selected position. We then replace each [MASK] with words predicted by BERT, resulting in an augmented instruction containing additional, contextually plausible tokens. This perturbation may introduce noise, redundancy, or shifts in meaning, challenging the model’s ability to extract the core task intention from a more verbose or distorted input.
@@ 4.1 Models
We experiment with two open-weight base LLMs in two sizes: Qwen-2.5 (7B and 72B) (yang2024qwen2); and Llama-3.1 (8B and 70B) (dubey2024llama).
@@ 4.2 Instruction Datasets
GPT4-Alpaca (peng2023instruction) is derived from Alpaca (alpaca), where the original examples are replaced with responses generated by GPT-4. Super-Natural Instruction (wang2022super) contains diverse tasks, including text classification and translation, with corresponding instructions. It is designed to evaluate the LLM abilities across a wide range of linguistic and functional contexts. Dolly (conover2023free) consists of instruction-following examples that reflect practical, real-world tasks like brainstorming and creative writing. The prompt-response pairs are high-quality and human-generated. Table 2 summarizes the number of samples in these datasets.
@@ 4.3 Implementation Details
We apply parameter-efficient fine-tuning methods for all experiments. Specifically, we use LoRA (hu2022lora) to fine-tune the 7B and 8B models, and QLoRA (dettmers2023qlora) for the larger 70B and 72B models. Each model is fine-tuned for one epoch on each dataset mixture to ensure consistency across experiments. All fine-tuning runs were performed on a single NVIDIA H100 GPU. Full details on the fine-tuning hyperparameters are provided in Appendix A.
@@ 4.4 Evaluation
Massive Multitask Language Understanding (MMLU; hendrycks2021measuring):
@@
Big-Bench Hard (BBH; Suzgun2022ChallengingBT):
A challenging subset of 23 tasks from the original BIG-Bench (srivastava2022beyond), aimed at evaluating advanced reasoning in language models. We assess both direct prompting and chain-of-thought (CoT) 10.5555/3600270.3602070, using official prompts with three in-context examples, and report average exact match across sub-tasks.
Grade School Math (GSM8K; Cobbe2021TrainingVT):
@@
Additionally, to analyze potential side effects of instruction-tuning on noisy instructions, such as changes in toxicity or misinformation generation, we evaluate the models on: (1) ToxiGen (hartvigsen2022toxigen) which measures the extent to which models generate toxic language and hate speech when explicitly prompted to do so across various demographic groups. We report the percentage of toxic outputs identified using a RoBERTa model (liu2020roberta) fine-tuned for toxicity detection, as described by hartvigsen2022toxigen; and (2) TruthfulQA proposed by lin2022truthfulqa which assesses how effectively models can refrain from generating known falsehoods caused by misconceptions or false beliefs, while still generating informative and useful content. We use two off-the-shelf task-specific judge models developed by AllenAI based on Llama-2 (7B) (touvron2023llama) for measuring truthfulness [^3] https://huggingface.co/allenai/truthfulqa-truth-judge-llama2-7B and informativeness, following the setup of Groeneveld2024OLMoAT. In both datasets, we evaluate using only the original, unaltered prompts to measure model toxicity and truthfulness.
@@ 5 Results
Prior work has shown that CoT prompting outperforms direct prompting on benchmarks like BBH and GSM8K NEURIPS2022_8bb0d291; 10.5555/3600270.3602070, and our results confirm this, especially under instruction perturbation. On BBH (CoT), models like Llama-70B fine-tuned with fully perturbed instructions achieve the highest scores of 79.0% and 76.4% when evaluated with 25% and 50% perturbed instructions, respectively, indicating that CoT prompting benefits from increased robustness introduced during training. While Qwen-72B performs best when fine-tuned with less perturbed instructions, its relatively weaker performance under direct prompting suggests that incorporating perturbation during fine-tuning still contributes to improved generalization and reasoning robustness.
@@
We observe that the impact of instruction-tuning seems to vary by task. For example, MMLU shows relatively modest improvement from instruction-tuning, regardless of whether or not perturbations are applied. In contrast, BBH appears to consistently benefit from instruction-tuning, especially for Llama models. This observation aligns with findings from sun-dredze-2025-amuro, who suggest that certain tasks are already well-represented in a model’s pre-training data, leaving limited room for further gains through instruction-tuning. On the other hand, tasks that are underrepresented or poorly learned during pre-training can potentially see more substantial improvements as the model acquires new task-specific capabilities during instruction-tuning.
@@ 7 Theoretical Grounding of Fine-Tuning on Noisy Instructions
Our findings suggest that incorporating perturbations during instruction-tuning may not only enhance model robustness to noisy or perturbed inputs but may also yield improvements on standard, unperturbed instructions. A plausible explanation for this effect is that noisy instruction-tuning serves as an implicit form of regularization bishop1995training, potentially encouraging models to move beyond reliance on superficial linguistic patterns. Exposure to a wide spectrum of instruction formulations, including those containing syntactic or semantic anomalies, may discourage overfitting to narrow or canonical phrasing. These perturbations effectively broaden the training distribution, functioning as a form of data augmentation pmlr-v97-dao19b; hernandez2018data; vaibhav-etal-2019-improving, and may thereby help LLMs to learn more robust and generalizable task representations. A particularly noteworthy observation is that models fine-tuned with 100% perturbed instructions often achieve high accuracy, even when evaluated on standard instructions. This may suggest that the perturbations could act not merely as noise, but as a potential source of useful inductive bias that enhances generalization across prompt formats.
@@ Acknowledgments
We thank Atsuki Yamaguchi and the anonymous reviewers for their invaluable feedback.
AA is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation grant EP/S023062/1. XT and NA are supported by the EPSRC [grant number EP/Y009800/1], through funding from Responsible AI UK (KP0016) as a Keystone project. We acknowledge the IT Services at the University of Sheffield for Supercomputing for the provision of HPC services.
@@ Appendix A Instruction-tuning Hyperparameters
For instruction-tuning and evaluation, we adopt the implementation from Open Instruct (wang2023far; ivison2023camels; ivison2024unpacking; Lambert2024TLU3P). Table 6 shows the hyperparameters used in our instruction fine-tuning experiments.